OpenStudy (anonymous):

How would you solve 2=square root of 3x-7

OpenStudy (shamil98):

√3x-7 = 2 you multiply both sides by ^2 to get rid of the square root so (√3x-7)^2 =2^2 3x - 7 = 4 then solve for x.

OpenStudy (anonymous):

Ok thanks but do you know what it means to graphically estimate the solution?

OpenStudy (shamil98):

no, sorry.

OpenStudy (anonymous):

ok thanks for the answer